# Weighted Point Configurations with Hyperuniformity: An Ecological Example and Models

### v1

@@ 3.2 Poisson point process and Ginibre point process
PPP is a completely random configuration of points [2, 5]. For each point $\bm{X}_{j}=(X_{j},Y_{j})$, the coordinates $X_{j}$ and $Y_{j}$ are chosen independently from a uniform distributions in $[0,L_{x}]$ and $[0,L_{y}]$, respectively. We repeat the random choosing of point $N$ times independently. Figure 3 shows an obtained sample for $L_{x}=1626$, $L_{y}=895$, and $N=9853$. The density is $\rho\fallingdotseq 6.771\times 10^{-3}$. The obtained PPP is denoted by $\Xi^{\rm PPP}$.
@@
$Z_{n}:=\begin{bmatrix}\frac{1}{\sqrt{2}}(X_{11}+iY_{11})&\cdots&\frac{1}{\sqrt{2}}(X_{1n}+iY_{1n})\\ \vdots&\ddots&\vdots\\ \frac{1}{\sqrt{2}}(X_{n1}+iY_{n1})&\cdots&\frac{1}{\sqrt{2}}(X_{nn}+iY_{nn})\end{bmatrix},$
@@
$\rho_{\mathrm{G}}L_{x}^{0}L_{y}^{0}=N\ \Longrightarrow\ L_{x}^{0}=\sqrt{\frac{N\pi}{\lambda}}.$ (3.7)
If we put $N=9853$ and $\lambda=L_{x}/L_{y}=1626/895$, (3.7) gives
@@
In our numerical calculation, we have used $Z_{n}$ with $n=2.2\times 10^{4}$. Hence we see $237.1<2\sqrt{n}\fallingdotseq 296$, and confirmed that the matrix size is large enough for our purpose. Then we dilate the obtained point configuration in the region $L_{x}^{0}\times L_{y}^{0}$ to that in $L_{x}\times L_{y}$; that is, the size of region is enlarged by linear factor $L_{x}/L_{x}^{0}\fallingdotseq 6.857$. The obtained sample is shown by Fig.3 (b). The expected density of dots is $\rho=(1/\pi)\times 6.857^{-2}\fallingdotseq 6.770\times 10^{-3}$. The exact number of points in Fig.3 (b) is 9862, which has a small deviation from the target number $N=9862$ due to a sample fluctuation and the finite-size effect of matrix. But this deviation is negligible in our scaling analysis mentioned below. The obtained GPP is denoted by $\Xi^{\rm GPP}$.
(a) Poisson point process $\Xi^{\rm PPP}$
(b) Ginibre point process $\Xi^{\rm GPP}$
Figure 3: Numerically obtained samples of PPP and GPP. Both are prepared in the region $L_{x}\times L_{y}=1626\times 895$ with density $\rho\fallingdotseq 6.77\times 10^{-3}$ matched with the digital data for bush 1 shown by Fig. 1.
@@ 3.3 Measurements of hyperuniformity by finite-size data
We consider a rectangular region $L_{x}\times L_{y}$, We fix the aspect ratio $\lambda=L_{y}/L_{x}$ and write it simply as $\Lambda_{L_{x}}$ putting only the size in the $x$-direction as a subscript. Let $m\in\mathbb{N}$ and we divide $\Lambda_{L_{x}}$ into $m\times m$ subregions. The lengths in the $x$- and $y$-directions are then given by
$\ell_{x}=L_{x}/m,\quad\ell_{y}=L_{y}/m.$
The obtained subregions are labeled by $j=1,2,\dots,m^{2}$ as
@@
First we consider an unmarked point process $\Xi(\Lambda_{L_{x}})$. Examples are the point configurations of center-of-mass coordinates of bushes in a desert $\Xi_{\Pi}$ for which the mass information is deleted, $\Xi^{\rm PPP}$, and $\Xi^{\rm GPP}$. For all $j=1,2,\dots,m^{2}$, the expectation of $\Xi(\Lambda_{\ell_{x}}^{j})$ is simply evaluated as $\langle\Xi(\Lambda_{\ell_{x}}^{j})\rangle\equiv\rho\ell_{x}\ell_{y}$, where $\rho$ is the density of points of the given sample $\Lambda_{L_{x}}$, $\rho=\Xi(\Lambda_{L_{x}})/(L_{x}L_{y})=N/(L_{x}L_{y})$. Then for each subregion $\Lambda_{\ell_{x}}^{j}$, $j=1,2,\dots,m^{2}$, we count the deviation of number of points observed in it from the expectation, and define the variance by the following arithmetic mean over $m^{2}$ subregions,
$\mathrm{Var}[\Xi(\Lambda_{\ell_{x}})]=\frac{1}{m^{2}}\sum_{j=1}^{m^{2}}\left(\Xi(\Lambda_{\ell_{x}}^{j})-\rho\ell_{x}\ell_{y}\right)^{2}.$ (3.8)
@@
$R^{\rm point}_{\ell_{x}}=\frac{\mathrm{Var}[\Xi_{\Pi}(\Lambda_{\ell_{x}})]}{\rho\ell_{x}\ell_{y}},\quad R^{\rm PPP}_{\ell_{x}}=\frac{\mathrm{Var}[\Xi^{\rm PPP}(\Lambda_{\ell_{x}})]}{\rho\ell_{x}\ell_{y}},\quad R^{\rm GPP}_{\ell_{x}}=\frac{\mathrm{Var}[\Xi^{\rm GPP}(\Lambda_{\ell_{x}})]}{\rho\ell_{x}\ell_{y}}.$ (3.9)
Next we consider an marked point process $\Pi(\Lambda_{L_{x}})$. Examples are the point configurations of center-of-mass coordinates of bushes in a desert $\Xi_{\Pi}$ weighted by masses. Instead of the density of point $\rho$, here we calculate the density of mass defined by
@@
For all $j=1,2,\dots,m^{2}$, the expectation of $\Pi(\Lambda_{\ell_{x}}^{j})$ is evaluated as $\langle\Xi(\Pi_{\ell_{x}}^{j})\rangle\equiv\rho^{\rm mass}\ell_{x}\ell_{y}$. Then the variance is calculated as
$\mathrm{Var}[\Pi(\Lambda_{\ell_{x}})]=\frac{1}{m^{2}}\sum_{j=1}^{m^{2}}\left(\Pi(\Lambda_{\ell_{x}}^{j}))-\rho^{\rm mass}\ell_{x}\ell_{y}\right)^{2}.$ (3.10)
@@
$R^{\rm mass}_{\ell_{x}}=\frac{\mathrm{Var}[\Pi(\Lambda_{\ell_{x}})]}{\rho^{\rm mass}\ell_{x}\ell_{y}}.$ (3.11)
@@ 3.4 Hyperuniformity of marked point processes of bushes in a desert in Argentina
(a) Log-log plots of $R_{\ell_{x}}$ versus $\ell_{x}$ for various unmarked and marked point processes. The bush configuration not taking into account of mass information does not show hyperuniformity similarly to PPP, while the bush configurations weighted by masses exhibit hyperuniformity similarly to GPP.
(b) Linear fitting to (3.12) were performed for the interval of $X\equiv\log\ell_{x}$ such that $2.8\leq X\leq 4.4$ for GPP, $2.8\leq X\leq 4.0$ for the marked point processes of bush 1-3 and 5, and $2.8\leq X\leq 3.8$ for bush 4. The fitting works well and the exponent $\alpha$ is evaluated as (3.13). Here the dots are shifted in the direction of $Y\equiv\log R$ by $-(n-1)$ for each bush n, $n=1,2,\dots,5$, respectively, to avoid overlaps of plots and fitting lines.
Figure 5: Numerical measurements of hyperuniformity using finite-size data
In Fig. 5 (a), we show dependence on $\log\ell_{x}$ of logarithms of $R^{\rm point}_{\ell_{x}}$ for the unmarked point configurations of the sample bush 1, $\log R^{\rm PPP}_{\ell_{x}}$ and $\log R^{\rm GPP}_{\ell x}$. For the totally five samples from the desert in Argentina, bush 1–5 in Table 1, $\log R^{\rm mass}_{\ell_{x}}$ are also plotted versus $\log\ell_{x}$, which are marked point processes in Fig. 5 (a). In the numerical measurements of variances by (3.8) and (3.10) in Section 3.3, the numbers of subregions $m$ become relatively small for large values of $\ell_{x}$, and hence the obtained plots are scattered. For this reason we only use the data with $\log\ell_{x}<4.5$.
With increment of the window size $\ell_{x}$, the number fluctuation $R^{\rm PPP}_{\ell_{x}}$ does not show any decay, while $R^{\rm GPP}_{\ell_{x}}$ shows a systematic decay. These plots suggest that $\Xi^{\rm PPP}$ has no hyperuniformity, but $\Xi^{\rm GPP}$ is in a hyperuniform state, as expected. We see that the behavior of $\log R^{\rm point}_{\ell_{x}}$ versus $\log\ell_{x}$ of unmarked point processes of bush 1 is very similar to that of PPP, as already observed in Fig. 5 (a) It implies that if the mass is not taken into account, any hyperuniformity can not be observed in bush configurations in the desert. If we consider the marked point processes, however, in which each point is weighted by mass, linear decay of $\log R^{\rm pass}_{\ell_{x}}$ with increment of $\ell_{x}$ is commonly observed as shown by the five data from the desert in Argentina, bush 1–5 (marked) in Fig. 5 (a).
@@
$R_{\ell_{x}}^{\rm mass}\simeq\ell_{x}^{-\alpha^{\rm Argentina}}\quad\mbox{with}\quad\alpha^{\rm Argentina}\fallingdotseq 0.65.$ (3.13)
@@ 3.5 Mass distributions of marked point processes of bushes
${\rm Prob}(M\in[m,m+dm))\simeq e^{-m/m_{0}}dm\quad\mbox{with}\quad m_{0}^{\rm Argentina}\fallingdotseq 77.$ (3.14)
@@ 4.1 Random thinning-coalescing processes
$\Lambda_{L_{x},r}:=\left\{(x,y)\in\mathbb{R}^{2}:-2r\leq x\leq L_{x}+2r,-2r\leq y\leq\lambda L_{x}+2r\right\}\supset\Lambda_{L_{x}}.$
@@
Choose a point $\bm{X}$ randomly from the set of points $\mathcal{P}$ which are included in $\Lambda_{L_{x}}$; $\bm{X}\in\mathcal{P}\cap\Lambda_{L_{x}}$.
Let ${A}$ be a non-negative random variable following the probability law (4.15) and set ${R:=\sqrt{A/\pi}}$. That is, $R$ is the radius of a disk whose area is equal to $A$. For a chosen point $\bm{X}$, define a subset of $\mathcal{P}$ by
$\mathcal{P}(\bm{X}):=\{\bm{Y}\in\mathcal{P}\backslash\left\{\bm{X}\right\}:d(\bm{X},\bm{Y})\leq R\}.$
Notice that $\mathcal{P}(\bm{X})$ is a collection of all points except $\bm{X}$, which are included in the disk centered at $\bm{X}$ with radius $R$.
Set
$\widetilde{\mathcal{P}}$ $=\mathcal{P}\backslash\mathcal{P}(\bm{X}),$ (4.16)
$\widetilde{M}(\bm{Z})$ =$=\begin{cases}\displaystyle{M(\bm{X})+\sum_{\bm{Y}:\bm{Y}\in\mathcal{P}(\bm{X})}M(\bm{Y})},&\mbox{if ${\bm{Z}=\bm{X}}$}\cr M(\bm{Z}),&\mbox{if ${\bm{Z}\in\widetilde{\mathcal{P}}\backslash\left\{\bm{X}\right\}}$}.\end{cases}$ (4.17)
The reduction of points (4.16) represents a random thinning [16, 17, 31]. That is, all points except $\bm{X}$ included in a disk centered at $\bm{X}$ with radius $R$ are deleted. Then the first line of (4.17) represents coalescing (aggregation, clustering) of masses of the deleted points into the chosen point $\bm{X}$ [8, 22, 25]. As reported in the following, we have iterated the transformation ${\mathcal{T}}$, where the times of iterations is written as $T$. We have found that even though the initial marked point process has no hyperuniformity, if we iterate the transformation $\mathcal{T}$ sufficiently many times, $T\gg 1$, then we can obtain non-trivial marked point process having hyperuniformity.

### v2

@@ 3.2 Poisson point process and Ginibre point process
Refer to caption Refer to caption
Figure 3: Numerically obtained samples of (a) Poisson point process (PPP) $\Xi^{\rm PPP}$ and (b) Ginibre point process (GPP) $\Xi^{\rm GPP}$. Both are prepared in the region $L_{x}\times L_{y}=1626\times 895$ with density $\rho\fallingdotseq 6.77\times 10^{-3}$ matched with the digital data for bush 1 shown by Fig. 1.
@@
PPP is a completely random configuration of points [1, 2]. For each point $\bm{X}_{j}=(X_{j},Y_{j})$, the coordinates $X_{j}$ and $Y_{j}$ are chosen independently from a uniform distributions in $[0,L_{x}]$ and $[0,L_{y}]$, respectively. We repeat the random choosing of point $N$ times independently. Figure 3 (a) shows an obtained sample for $L_{x}=1626$, $L_{y}=895$, and $N=9853$. The density is $\rho\fallingdotseq 6.771\times 10^{-3}$. The obtained PPP is denoted by $\Xi^{\rm PPP}$.
@@
$Z_{n}=\begin{bmatrix}\frac{1}{\sqrt{2}}(X_{11}+iY_{11})&\cdots&\frac{1}{\sqrt{2}}(X_{1n}+iY_{1n})\\ \vdots&\ddots&\vdots\\ \frac{1}{\sqrt{2}}(X_{n1}+iY_{n1})&\cdots&\frac{1}{\sqrt{2}}(X_{nn}+iY_{nn})\end{bmatrix},$
@@
$\rho_{\mathrm{G}}L_{x}^{0}L_{y}^{0}=N\ \Longrightarrow\ L_{x}^{0}=\sqrt{\frac{N\pi}{\lambda}}.$ (7)
If we put $N=9853$ and $\lambda=L_{y}/L_{x}=895/1626$, (7) gives
@@
In our numerical calculation, we have used $Z_{n}$ with $n=2.2\times 10^{4}$. Hence we see $237.1<2\sqrt{n}\fallingdotseq 296$, and confirmed that the matrix size is large enough for our purpose. Then we dilate the obtained point configuration in the region $L_{x}^{0}\times L_{y}^{0}$ to that in $L_{x}\times L_{y}$; that is, the size of region is enlarged by linear factor $L_{x}/L_{x}^{0}\fallingdotseq 6.857$. The obtained sample is shown by Fig.3 (b). The expected density of dots is $\rho=(1/\pi)\times 6.857^{-2}\fallingdotseq 6.770\times 10^{-3}$. The exact number of points in Fig.3 (b) is 9862, which has a small deviation from the target number $N=9853$ due to a sample fluctuation and the finite-size effect of matrix. But this deviation is negligible in our scaling analysis mentioned below. The obtained GPP is denoted by $\Xi^{\rm GPP}$.
@@ 3.3 Measurements of hyperuniformity by finite-size data
We consider a rectangular region $L_{x}\times L_{y}$. We fix the aspect ratio $\lambda=L_{y}/L_{x}$ and write it simply as $\Lambda_{L_{x}}$ putting only the size in the $x$-direction as a subscript. Let $m\in\mathbb{N}$ and we divide $\Lambda_{L_{x}}$ into $m\times m$ subregions. The lengths in the $x$- and $y$-directions are then given by
$\ell_{x}=L_{x}/m\quad\ell_{y}=L_{y}/m.$
See Fig.4. The obtained subregions are labeled by $j=1,2,\dots,m^{2}$ as
@@
First we consider an unmarked point process $\Xi(\Lambda_{L_{x}})$. Examples are the point configurations of center-of-mass coordinates of bushes in a desert $\Xi_{\Pi}$ for which the mass information is deleted, $\Xi^{\rm PPP}$, and $\Xi^{\rm GPP}$. For all $j=1,2,\dots,m^{2}$, the expectation of $\Xi(\Lambda_{\ell_{x}}^{j})$ is simply evaluated as $\langle\Xi(\Lambda_{\ell_{x}}^{j})\rangle=\rho\ell_{x}\ell_{y}$, where $\rho$ is the density of points of the given sample $\Lambda_{L_{x}}$, $\rho=\Xi(\Lambda_{L_{x}})/(L_{x}L_{y})=N/(L_{x}L_{y})$. Then for each subregion $\Lambda_{\ell_{x}}^{j}$, $j=1,2,\dots,m^{2}$, we count the deviation of number of points observed in it from the expectation, and define the variance by the following arithmetic mean over $m^{2}$ subregions,
$\mathrm{Var}[\Xi(\Lambda_{\ell_{x}})]=\frac{1}{m^{2}}\sum_{j=1}^{m^{2}}\left(\Xi(\Lambda_{\ell_{x}}^{j})-\rho\ell_{x}\ell_{y}\right)^{2}.$ (8)
@@
$R^{\rm point}_{\ell_{x}}=\frac{\mathrm{Var}[\Xi_{\Pi}(\Lambda_{\ell_{x}})]}{\rho\ell_{x}\ell_{y}},\quad R^{\rm PPP}_{\ell_{x}}=\frac{\mathrm{Var}[\Xi^{\rm PPP}(\Lambda_{\ell_{x}})]}{\rho\ell_{x}\ell_{y}},\quad R^{\rm GPP}_{\ell_{x}}=\frac{\mathrm{Var}[\Xi^{\rm GPP}(\Lambda_{\ell_{x}})]}{\rho\ell_{x}\ell_{y}}.$ (9)
Next we consider a marked point process $\Pi(\Lambda_{L_{x}})$. Examples are the point configurations of center-of-mass coordinates of bushes in a desert $\Xi_{\Pi}$ weighted by masses. Instead of the density of point $\rho$, here we calculate the density of mass defined by
@@
For all $j=1,2,\dots,m^{2}$, the expectation of $\Pi(\Lambda_{\ell_{x}}^{j})$ is evaluated as $\langle\Pi(\Lambda_{\ell_{x}}^{j})\rangle=\rho^{\rm mass}\ell_{x}\ell_{y}$. Then the variance is calculated as
$\mathrm{Var}[\Pi(\Lambda_{\ell_{x}})]=\frac{1}{m^{2}}\sum_{j=1}^{m^{2}}\left(\Pi(\Lambda_{\ell_{x}}^{j}))-\rho^{\rm mass}\ell_{x}\ell_{y}\right)^{2}.$ (10)
@@
$R^{\rm mass}_{\ell_{x}}=\frac{\mathrm{Var}[\Pi(\Lambda_{\ell_{x}})]}{\rho^{\rm mass}\ell_{x}\ell_{y}}.$ (11)
@@ 3.4 Hyperuniformity of marked point processes of bushes in a desert in Argentina
Figure 5: Numerical measurements of hyperuniformity using finite-size data. (a) Log-log plots of $R_{\ell_{x}}$ versus $\ell_{x}$ for various unmarked and marked point processes. The bush configuration not taking into account of mass information does not show hyperuniformity similarly to PPP, while the bush configurations weighted by masses exhibit hyperuniformity similarly to GPP. (b) Linear fitting to (12) was performed for the interval of $X\equiv\log\ell_{x}$ such that $2.8\leq X\leq 4.4$ for GPP, $2.8\leq X\leq 4.0$ for the marked point processes of bush 1-3 and 5, and $2.8\leq X\leq 3.8$ for bush 4. The fitting works well and the exponent $\alpha$ is evaluated as (13). Here the dots are shifted in the direction of $Y\equiv\log R$ by $-(n-1)$ for each bush n, $n=1,2,\dots,5$, respectively, to avoid overlaps of plots and fitting lines.
In Fig. 5 (a), we show dependence on $\log\ell_{x}$ of logarithms of $R^{\rm point}_{\ell_{x}}$ for the unmarked point configurations of the sample bush 1, $\log R^{\rm PPP}_{\ell_{x}}$ and $\log R^{\rm GPP}_{\ell x}$. For the totally five samples from the desert in Argentina, bush 1–5 in Table 1, $\log R^{\rm mass}_{\ell_{x}}$ are also plotted versus $\log\ell_{x}$ in Fig. 5 (a), which are for the marked point processes. In the numerical measurements of variances by (8) and (10) in Section 3.3, the numbers of subregions $m$ become relatively small for large values of $\ell_{x}$, and hence the obtained plots are scattered. For this reason we only use the data with $\log\ell_{x}<4.5$.
With increment of the window size $\ell_{x}$, the number fluctuation $R^{\rm PPP}_{\ell_{x}}$ does not show any decay, while $R^{\rm GPP}_{\ell_{x}}$ shows a systematic decay. These plots suggest that $\Xi^{\rm PPP}$ has no hyperuniformity, but $\Xi^{\rm GPP}$ is in a hyperuniform state, as expected. We see that the behavior of $\log R^{\rm point}_{\ell_{x}}$ versus $\log\ell_{x}$ of unmarked point processes of bush 1 is very similar to that of PPP, as already observed in Fig. 1. It implies that if the mass is not taken into account, any hyperuniformity can not be observed in bush configurations in the desert. If we consider the marked point processes, however, in which each point is weighted by mass, linear decay of $\log R^{\rm pass}_{\ell_{x}}$ with increment of $\log\ell_{x}$ is commonly observed as shown by the five data from the desert in Argentina, bush 1–5 (marked) in Fig. 5 (a).
@@
$R_{\ell_{x}}^{\rm mass}\simeq\ell_{x}^{-\alpha^{\rm Argentina}}\quad\mbox{with}\quad\alpha^{\rm Argentina}\fallingdotseq 0.65.$ (13)
@@ 3.5 Mass distributions of marked point processes of bushes
${\rm Prob}(M\in[\mathsf{m},\mathsf{m}+d\mathsf{m}))\simeq e^{-\mathsf{m}/\mathsf{m}_{0}}d\mathsf{m}\quad\mbox{with}\quad\mathsf{m}_{0}=\mathsf{m}_{0}^{\rm Argentina}\fallingdotseq 77.$ (14)
@@ 4.1 Random thinning-coalescing processes
$\Lambda_{L_{x},r}=\left\{(x,y)\in\mathbb{R}^{2}:-2r\leq x\leq L_{x}+2r,-2r\leq y\leq\lambda L_{x}+2r\right\}\supset\Lambda_{L_{x}}.$
@@
(i)  Choose a point $\bm{X}$ randomly from the set of points $\mathcal{P}$ which are included in $\Lambda_{L_{x}}$; $\bm{X}\in\mathcal{P}\cap\Lambda_{L_{x}}$.
(ii)  Let $A$ be a non-negative random variable following the probability law (15) and set $R\equiv\sqrt{A/\pi}$. That is, $R$ is the radius of a disk whose area is equal to $A$. For a chosen point $\bm{X}$, define a subset of $\mathcal{P}$ by
$\mathcal{P}(\bm{X})=\{\bm{Y}\in\mathcal{P}\backslash\left\{\bm{X}\right\}:d(\bm{X},\bm{Y})\leq R\},$
where $A\setminus B$ means subtraction of a set $B$ from a set $A$. Notice that $\mathcal{P}(\bm{X})$ is a collection of all points except $\bm{X}$, which are included in the disk centered at $\bm{X}$ with radius $R$.
(iii)  Set
$\widetilde{\mathcal{P}}$ $=\mathcal{P}\backslash\mathcal{P}(\bm{X}),$ (16)
$\widetilde{M}(\bm{Z})$ =$=\begin{cases}\displaystyle{M(\bm{X})+\sum_{\bm{Y}:\bm{Y}\in\mathcal{P}(\bm{X})}M(\bm{Y})},&\mbox{if ${\bm{Z}=\bm{X}}$}\cr M(\bm{Z}),&\mbox{if ${\bm{Z}\in\widetilde{\mathcal{P}}\backslash\left\{\bm{X}\right\}}$}.\end{cases}$ (17)
The reduction of points (16) represents a random thinning [21, 22, 23]. That is, all points except $\bm{X}$ included in a disk centered at $\bm{X}$ with radius $R$ are deleted. Then the first line of (17) represents coalescing (aggregation, clustering) of masses of the deleted points into the chosen point $\bm{X}$ [24, 25, 26].
As reported in the following, we have iterated the transformation ${\mathcal{T}}$. We have found that even though the initial marked point process has no hyperuniformity, if we iterate the transformation $\mathcal{T}$ sufficiently many times, then we can obtain non-trivial marked point process having hyperuniformity. In the practical simulation, iteration of $\mathcal{T}$ is performed as follows. First we fix a rate $p\in(0,1)$. Assume that in a given marked point process $\Pi=\sum_{\bm{X}:\bm{X}\in\mathcal{P}}M(\bm{X})\delta_{\bm{X}}$, the total number of points is $n=\sharp(\mathcal{P}\cap\Lambda_{L_{x}})=\Xi_{\Pi}(\Lambda_{L_{x}})$. We choose $pn$ points from $\mathcal{P}\cap\Lambda_{L_{x}}$ uniformly at random and prepare an ordered list of them as candidates for the central point of a disk in the procedures (ii) and (iii) in the above algorithm. We assign the first point in the list as $\bm{X}$ in (i) and then perform (ii) and (iii). In the thinning (16), if some of the points in the list are deleted, then they are deleted also in the list. We eliminate also the used point from the list. Next in the updated list the first point is assigned as $\bm{X}$ in (i) and using it as a center of a disk, the procedures (ii) and (iii) are performed. By definition, the number of points in the list decreases monotonically. When the list becomes empty, we think that one iteration-term is complete and increase the number of iteration-terms $T$ by unity; $T\to T+1$. For the updated marked point process $\widetilde{\Pi}=\sum_{\bm{X}:\bm{X}\in\widetilde{\mathcal{P}}}\widetilde{M}(\bm{X})\delta_{\bm{X}}$, where the total number of points is now given by $\widetilde{n}=\sharp(\widetilde{\mathcal{P}}\cap\Lambda_{L_{x}})$, we choose $p\widetilde{n}$ points randomly from $\widetilde{\mathcal{P}}\cap\Lambda_{L_{x}}$ and make a new ordered list of candidates for the center of a disk. We repeat the iteration-terms of $\mathcal{T}$. By the thinning (16), the total number of point $n$ is decreasing, and hence the computational time for one iteration-term decreases as $T$ increases, since we fix $p\in(0,1)$. In the following, we report the results in the case $p=0.1$.